# Inertial Coordination Games

### v2

@@ A.2  Theorem 1 and Corollary 2
[…]
@@
[…]
@@
[…]
Fix any $q>2$, $C>0$, $T\geq 1$, and $\Sigma$. Fix any $q^{\prime}\in(2,q)$, and choose $T^{\prime}\geq T$ large enough such that
$C(T^{\prime})^{q-q^{\prime}}\geq\underline{C}(q^{\prime},\lambda_{0}),$
where $\underline{C}(q^{\prime},\lambda_{0})$ is the constant from Lemma 8. Then, the following implications hold:
@@
[…]
@@
Part (ii). If $\sigma_{1}:=\sigma$ and $\sigma_{t}:=\eta_{t-1}$ for all $t\geq 2$, then
$\eta_{t}^{-2}=\sigma^{-2}2^{t-1}$
Fix $q=3$ and any $\lambda_{0}$, and let $\underline{C}$ be the constant from Lemma 8 when $q=2$. Choose $\underline{\sigma}>0$ small enough such that
$\underline{\sigma}^{-2}2^{t-1}\geq\underline{C}t^{3}\quad\forall\ t\geq 1,$
and note that $\eta_{t}^{-2}\geq\underline{\sigma}^{-2}2^{t-1}$ for all $\sigma<\underline{\sigma}$. The conclusion then follows from Lemma 8. ∎
@@ A.4  Proposition 2
By taking $t\to\infty$ for ($\mu^{*}$-LoM), we see that
$\mu_{\infty}^{*}+\Phi(\gamma_{\infty})=c$
and hence
@@
Step 2. To complete the proof of Proposition 2, we show that there exists a noise process $\Sigma=(\sigma_{t})_{t}$ that implements the sequence $(A_{t})_{t}$ constructed above. First, we claim that it suffices to find a non-negative, weakly decreasing sequence $(\eta_{t}^{2})_{t}$ of posterior variances. This is true because given any such $(\eta_{t}^{2})_{t}$, the sequence $(\eta_{t}^{-2})_{t}$ is non-negative and weakly increasing. Thus, $(\eta_{t}^{2})_{t}$ is implemented by the noise process
$\sigma_{1}^{-2}:=\eta_{1}^{-2},\quad\sigma_{t}^{-2}=\eta_{t}^{-2}-\eta_{t-1}^{-2}\quad\forall\ t\geq 2$
Thus, it suffices to find a non-negative and weakly decreasing sequence $(\eta_{t}^{2})_{t\geq 1}$ such that
$A_{t}^{2}=\eta_{t}^{2}+\eta_{t-1}^{2}\quad\forall t\geq 2$
Note that any choice of $\eta_{1}^{2}$ pins down a candidate solution $(\eta_{t}^{2})_{t\geq 1}$ by $\eta_{t+1}^{2}=A_{t+1}^{2}-\eta_{t}^{2}$. We therefore proceed by stating the implied constraints on $\eta_{1}^{2}$.
@@
For every $t\geq 2$, we have the constraint
$0\leq\eta_{t+1}^{2}\leq\eta_{t}^{2}$
We can split this set of constraints into two cases.
@@ A.5  Proposition 3
First, we show that $T(\sigma)\to\infty$ as $\sigma\downarrow 0$. Notice that $T(\sigma)$ is weakly decreasing in $\sigma$ because, by the proof of Proposition 1(ii), for any $\sigma^{\prime}>\sigma$, $\theta\geq\mu^{*}_{T(\sigma^{\prime})}(\sigma^{\prime})>\mu^{*}_{T(\sigma^{\prime})}(\sigma)$. Hence, $T(\sigma)\geq T(\sigma^{\prime})$ by the definition of $T(\sigma)$. Hence, $T(\sigma)$ either converges to some constant or diverges to $\infty$ as $\sigma\downarrow 0$. Suppose for a contradiction that $\lim_{\sigma\downarrow 0}T(\sigma)=T$ holds for some $T>0$. Since $T(\sigma)\uparrow T$ as $\sigma\downarrow 0$, for any $\sigma$, we have
$\mu^{*}_{T+1}(\sigma)>\mu^{*}_{T(\sigma)+1}(\sigma)>\theta,$
which implies $\lim_{\sigma\downarrow 0}\mu^{*}_{T+1}(\sigma)\geq\theta$. This contradicts $\lim_{\sigma\downarrow 0}\mu^{*}_{T+1}(\sigma)=c-\lambda_{0}$.
@@
Finally, fix any $\beta\in(1,\overline{\beta})$, where
$\overline{\beta}:=c+(1/2)-\theta-\delta\quad\text{and}\quad\delta=(c-(1/2)-\theta)/2$
and note that, for any $t\in[\beta T(\sigma),\infty)$,
@@
Hence, for any $\epsilon$, there exists $\overline{\sigma}>0$ such that $|\lambda_{t+1}-\lambda_{t}|<\epsilon$ holds for all $t\geq 1$ and $\sigma>\overline{\sigma}$. Note that this argument does not depend on the value of $\theta$ as long as $|\theta|<\infty$ holds.
Appendix B Small Lags
@@
$(\eta_{t}^{(n)})^{2}=\left(\sum_{k\in\mathcal{T}^{(n)},k\leq t}\frac{1}{(\sigma_{t}^{(n)})^{2}}\right)^{-1}.$
@@
    Consider the model with contemporaneous incentives, in which $U(1,t)=\theta+\lambda_{t}$. Then, the unique symmetric threshold strategy is $\mu_{t}^{*}=c-(1/2)$.
@@
[…]
@@
Part 1. Fix any $t\in\mathcal{T}$. Suppose that agents follow the same threshold strategy; that is, agent $i$ chooses action $1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$, and this is common knowledge. Agent $i$’s expected payoff of playing $1$ is
$\mathbb{E}_{it}[\theta+\lambda_{t}(\theta)]=\mu_{it}+\mathbb{P}_{it}(\mu_{jt}\geq\mu_{t}^{*})=\mu_{it}+\Phi\left(\frac{\mu_{it}-\mu_{t}^{*}}{\sqrt{2}\eta_{t}}\right).$
Since she is indifferent between playing 00 and $1$ when $\mu_{it}=\mu_{t}^{*}$, it must be that
$\mu_{t}^{*}+\Phi\left(\frac{\mu_{t}^{*}-\mu_{t}^{*}}{\sqrt{2}\eta_{t}}\right)=c\iff\mu_{t}^{*}=c-\frac{1}{2}.$
@@
and $x(\theta)=\min\Big{\{}|x_{1}(\theta)|,|x_{2}(\theta)|\Big{\}}$. Observe that
@@
$\leq\theta\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2}\pi}\frac{\eta_{t-1/n}^{(n)}\mu_{t}^{(n)*}-\eta_{t}^{(n)}\mu_{t-1/n}^{(n)*}}{\eta_{t}^{(n)}\eta_{t-1/n}^{(n)}}$ ($\varphi(\cdot)\leq(\sqrt{2}\pi)^{-1}$, $\eta_{t-1/n}^{(n)}\geq\eta_{t}^{(n)}$, and $\mu_{t}^{(n)*}\geq\mu_{t-1/n}^{(n)*}$)
$=\theta\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}}{\sqrt{2}\pi\eta_{t}^{(n)}}+\frac{\eta_{t-1/n}^{(n)}-\eta_{t}^{(n)}}{\sqrt{2}\pi\eta_{t}^{(n)}\eta_{t-1/n}^{(n)}}\mu_{t-1/n}^{(n)*}$
$\leq\theta\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2}\pi\eta_{t}^{(n)}(nt-1)}$ ($\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}\leq(nt-1)^{-1}$ by (3))
$\leq\max\left\{|\mu_{t}^{(n)*}|+\eta_{t}^{(n)},|\mu_{t-1/n}^{(n)*}|+\eta_{t-1/n}^{(n)}\right\}\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2}\pi\eta_{t}^{(n)}(nt-1)}$
$:=\epsilon(n,t),$
@@
$\theta\varphi(|x_{1}(\theta)|)$ $=\left(\mu_{t}^{(n)*}+\eta_{t}^{(n)}x_{1}(\theta)\right)\varphi(|x_{1}(\theta)|)$
$\leq\left(|\mu_{t}^{(n)*}|+\eta_{t}^{(n)}x_{1}(\theta)\right)\varphi(|x_{1}(\theta)|)$
$\leq|\mu_{t}^{(n)*}|+\eta_{t}^{(n)}$ ($a\varphi(|a|)\leq\varphi(1)\leq 1$)
and $\theta\varphi(|x_{2}(\theta)|)\leq|\mu_{t-1/n}^{(n)*}|+\eta_{t-1/n}^{(n)}$ by analogous arguments.
@@
Hence, the time-$t$ strategy “$a_{it}=1$ iff $\mu_{it}\geq\mu_{t}^{(n)*}$” is an $\epsilon(n,t)$-equilibrium at time $t$, since we may bound the time-$t$ expected payoff of using this strategy in the contemporaneous game uniformly across players’ time-$t$ information sets $\mu_{it}$: if $\mu_{it}\geq\mu_{t}^{(n)*}$,
@@
Appendix C Signals about past play
@@
Now suppose that the statement holds for $t-1$. By assumption, we have $\Phi^{-1}(\lambda_{t-1})=(\theta-\mu_{t-1}^{*})/\eta_{t-1}$, which implies
@@
ONLINE APPENDIX TO ‘INERTIAL COORDINATION GAMES’
Kei Uzui [^2] Department of Economics, MIT, kuzui@mit.edu
FOR ONLINE PUBLICATION ONLY

### v3

@@ A.2 Theorem 1 and Corollary 2
$\eta_{t}^{-2}\geq\underline{C}t^{q}\quad\forall t\geq T\implies\lambda_{\infty}(\theta)=NRD(\theta)\quad\text{p.m.}$
@@
$L:=\frac{2^{-q}}{2^{-q}+1},\ \ \ r:=\frac{q}{2}>1,\ \ \ S:=\sum_{t=2}^{\infty}t^{-r}\leq\int_{t=1}^{t=\infty}t^{-r}\ dt=\frac{1}{r-1}<\infty,$
@@
$\eta_{t}^{-2}=\Omega(t^{q})\implies\lambda_{\infty}(\theta)=NRD(\theta)\quad\text{p.m.}$
Fix any $q>2$, $C>0$, $T\geq 1$, and $\Sigma$. Fix any $q^{\prime}\in(2,q)$, and choose $T^{\prime}\geq T$ large enough such that $C(T^{\prime})^{q-q^{\prime}}\geq\underline{C}(q^{\prime},\lambda_{0}),$ where $\underline{C}(q^{\prime},\lambda_{0})$ is the constant from Lemma 8. Then, the following implications hold:
@@
$\implies\lambda_{\infty}(\theta)=NRD(\theta)\quad\text{p.m.}$ (From Lemma 8)
@@
Part (ii). If $\sigma_{1}:=\sigma$ and $\sigma_{t}:=\eta_{t-1}$ for all $t\geq 2$, then $\eta_{t}^{-2}=\sigma^{-2}2^{t-1}$ holds. Fix $q=3$ and any $\lambda_{0}$, and let $\underline{C}$ be the constant from Lemma 8 when $q=2$. Choose $\underline{\sigma}>0$ small enough such that $\underline{\sigma}^{-2}2^{t-1}\geq\underline{C}t^{3}\quad\forall\ t\geq 1,$ and note that $\eta_{t}^{-2}\geq\underline{\sigma}^{-2}2^{t-1}$ for all $\sigma<\underline{\sigma}$. The conclusion then follows from Lemma 8. ∎
@@ A.4 Proposition 2
By taking $t\to\infty$ for ($\mu^{*}$-LoM), we see that $\mu_{\infty}^{*}+\Phi(\gamma_{\infty})=c$ and hence
@@
Step 2. To complete the proof of Proposition 2, we show that there exists a noise process $\Sigma=(\sigma_{t})_{t}$ that implements the sequence $(A_{t})_{t}$ constructed above. First, we claim that it suffices to find a non-negative, weakly decreasing sequence $(\eta_{t}^{2})_{t}$ of posterior variances. This is true because given any such $(\eta_{t}^{2})_{t}$, the sequence $(\eta_{t}^{-2})_{t}$ is non-negative and weakly increasing. Thus, $(\eta_{t}^{2})_{t}$ is implemented by the noise process $\sigma_{1}^{-2}:=\eta_{1}^{-2},$ and $\sigma_{t}^{-2}=\eta_{t}^{-2}-\eta_{t-1}^{-2}$ for $t\geq 2$. Thus, it suffices to find a non-negative and weakly decreasing sequence $(\eta_{t}^{2})_{t\geq 1}$ such that $A_{t}^{2}=\eta_{t}^{2}+\eta_{t-1}^{2}$ for $t\geq 2$. Note that any choice of $\eta_{1}^{2}$ pins down a candidate solution $(\eta_{t}^{2})_{t\geq 1}$ by $\eta_{t+1}^{2}=A_{t+1}^{2}-\eta_{t}^{2}$. We therefore proceed by stating the implied constraints on $\eta_{1}^{2}$.
@@
For every $t\geq 2$, we have the constraint $0\leq\eta_{t+1}^{2}\leq\eta_{t}^{2}.$ We can split this set of constraints into two cases.
@@ A.5 Proposition 3
First, we show that $T(\sigma)\to\infty$ as $\sigma\downarrow 0$. Notice that $T(\sigma)$ is weakly decreasing in $\sigma$ because, by the proof of Proposition 1(ii), for any $\sigma^{\prime}>\sigma$, $\theta\geq\mu^{*}_{T(\sigma^{\prime})}(\sigma^{\prime})>\mu^{*}_{T(\sigma^{\prime})}(\sigma)$. Hence, $T(\sigma)\geq T(\sigma^{\prime})$ by the definition of $T(\sigma)$. Hence, $T(\sigma)$ either converges to some constant or diverges to $\infty$ as $\sigma\downarrow 0$. Suppose for a contradiction that $\lim_{\sigma\downarrow 0}T(\sigma)=T$ holds for some $T>0$. Since $T(\sigma)\uparrow T$ as $\sigma\downarrow 0$, for any $\sigma$, we have $\mu^{*}_{T+1}(\sigma)>\mu^{*}_{T(\sigma)+1}(\sigma)>\theta,$ which implies $\lim_{\sigma\downarrow 0}\mu^{*}_{T+1}(\sigma)\geq\theta$. This contradicts $\lim_{\sigma\downarrow 0}\mu^{*}_{T+1}(\sigma)=c-\lambda_{0}$.
@@
Finally, fix any $\beta\in(1,\overline{\beta})$, where $\overline{\beta}:=c+(1/2)-\theta-\delta\quad\text{and}\quad\delta=(c-(1/2)-\theta)/2$, and note that, for any $t\in[\beta T(\sigma),\infty)$,
@@
Hence, for any $\epsilon$, there exists $\overline{\sigma}>0$ such that $|\lambda_{t+1}-\lambda_{t}|<\epsilon$ holds for all $t\geq 1$ and $\sigma>\overline{\sigma}$. Note that this argument does not depend on the value of $\theta$ as long as $|\theta|<\infty$ holds. ∎
APPENDIX TO ‘INERTIAL COORDINATION GAMES’
Kei Uzui [^2] Department of Economics, MIT, kuzui@mit.edu
FOR ONLINE PUBLICATION ONLY
Appendix II Small Lags
@@
$(\eta_{t}^{(n)})^{2}=\left(\sum_{k\in\mathcal{T}^{(n)},k\leq t}\frac{1}{(\sigma_{k}^{(n)})^{2}}\right)^{-1}.$
Note that for $t\in\mathbb{Z}$ and letting $s(k)=\text{ceil}(k)$, we may write:
$(\eta_{t}^{(n)})^{2}=\left(\sum_{k\in\mathcal{T}^{(n)},k\leq t}\frac{1}{n\sigma_{s(k)}^{2}}\right)^{-1}$
$=n\left(\sum_{k\in\mathcal{T}^{(n)},k\leq t}\frac{1}{\sigma_{s(k)}^{2}}\right)^{-1}$
$=n\left(n\frac{1}{\sigma_{1}^{2}}\cdots+n\frac{1}{\sigma_{t}^{2}}\right)^{-1}=\eta_{t}^{2}$
@@
    Consider the model with contemporaneous incentives, in which $U(1,t)=\theta+\lambda_{t}$. Then, the unique strategy that survives iterated deletion of strictly dominated strategies (IDSDS) is the symmetric cutoff strategy $\mu_{t}^{*}=c-(1/2)$, that is, everyone plays action $1$ if and only if her posterior mean is greater than c(1/2)c-(1/2).
@@
      $\eta^{-2}_{t}=O(t^{2-\epsilon})\implies\lambda_{\infty}^{(n)}(\theta)=RD(\theta)\quad\text{a.e.}$
@@
Part 1. Fix any $t\in\mathcal{T}$. Define $b(x)$ as the unique value of $\mu$ solving the following equation:
$\mu+\Phi\left(\frac{\mu-x}{\sqrt{2}\eta_{t}}\right)=c.$
We show by induction that a strategy $s^{k}$ surviving $k$ rounds of IDSDS satisfies
$s^{k}(\mu_{it})=\begin{cases}1&\text{if $\mu_{it}>b^{k-1}(c)$}\\ 0&\text{if $\mu_{it}<b^{k-1}(c-1)$},\end{cases}$
where $\mu_{it}$ denotes $i$’s posterior mean at time $t$, and $b^{0}(x)=x$ holds. First, the statement holds for $k=1$ because it is strictly dominant to play 11 (0) when $\mu_{it}\in(c,+\infty)$ ($\mu_{it}\in(-\infty,c-1)$). Next suppose that the statement is true for $k$. Since agents know that others play $1$ whenever their posterior mean is greater than $b^{k-1}(c)$, agent $i$’s expected payoff of playing $1$ is at least
$\mu_{it}+\mathbb{P}_{it}(\mu_{jt}>b^{k-1}(c))=\mu_{it}+\Phi\left(\frac{\mu_{it}-b^{k-1}(c)}{\sqrt{2}\eta_{t}}\right).$
This implies that any strategy surviving $k$ rounds such that $s^{k}(\mu_{it})=0$ for $\mu_{it}\in(b(b^{k-1}(c)),b^{k-1}(c)]$ is eliminated in the $(k+1)$-th round. By an analogous argument, $s^{k}(\mu_{it})=1$ for $\mu_{it}\in[b^{k-1}(c-1),b(b^{k-1}(c-1)))$ is eliminated. Hence, the statement is true for $k+1$. Finally, since $\lim_{k\to\infty}b^{k}(c)=\lim_{k\to\infty}b^{k}(c-1)=c-(1/2)$ holds, a cutoff strategy with threshold c(1/2)c-(1/2) is the unique strategy surviving IDSDS.
@@
and $x(\theta)=\min\Big{\{}|x_{1}(\theta)|,|x_{2}(\theta)|\Big{\}}$. Let $\overline{c}=\max\{|c|,|c-1|\}$ and observe that
@@
$\leq|\theta|\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2\pi}}\frac{|\eta_{t-1/n}^{(n)}\mu_{t}^{(n)*}-\eta_{t}^{(n)}\mu_{t-1/n}^{(n)*}|}{\eta_{t}^{(n)}\eta_{t-1/n}^{(n)}}$ ($\varphi(\cdot)\leq(\sqrt{2\pi})^{-1}$)
$\leq|\theta|\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\left|\frac{\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}}{\sqrt{2\pi}\eta_{t}^{(n)}}\right|+\left|\frac{\eta_{t-1/n}^{(n)}-\eta_{t}^{(n)}}{\sqrt{2\pi}\eta_{t}^{(n)}\eta_{t-1/n}^{(n)}}\mu_{t-1/n}^{(n)*}\right|$
$\leq|\theta|\varphi(x(\theta))\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2\pi}\eta_{t}^{(n)}(nt-1)}+\frac{\eta_{t-1/n}^{(n)}-\eta_{t}^{(n)}}{\sqrt{2\pi}\eta_{t}^{(n)}\eta_{t-1/n}^{(n)}}|\mu_{t-1/n}^{(n)*}|$ ($\mu_{t}^{(n)*}-\mu_{t-1/n}^{(n)*}\leq(nt-1)^{-1}$ by (3))
$=\left(|\theta|\varphi(x(\theta))+\frac{1}{\sqrt{2\pi}}|\mu_{t-1/n}^{(n)*}|\right)\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2\pi}\eta_{t}^{(n)}(nt-1)}$
$\leq\max\left\{|\mu_{t}^{(n)*}|+\eta_{t}^{(n)}+\frac{\overline{c}}{\sqrt{2\pi}},|\mu_{t-1/n}^{(n)*}|+\eta_{t-1/n}^{(n)}+\frac{\overline{c}}{\sqrt{2\pi}}\right\}\left(\frac{1}{\eta_{t}^{(n)}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)$
$+\frac{1}{\sqrt{2\pi}\eta_{t}^{(n)}(nt-1)}$
$\leq\left(2\overline{c}+\eta_{1}\right)\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1/n}^{(n)}}\right)+\frac{1}{\sqrt{2\pi}\eta_{t}(nt-1)}:=\epsilon(n,t),$
@@
$|\theta|\varphi(|x_{1}(\theta)|)$ $=\left|\mu_{t}^{(n)*}+\eta_{t}^{(n)}x_{1}(\theta)\right|\varphi(|x_{1}(\theta)|)$
$\leq\left(|\mu_{t}^{(n)*}|+\eta_{t}^{(n)}|x_{1}(\theta)|\right)\varphi(|x_{1}(\theta)|)$
$\leq|\mu_{t}^{(n)*}|+\eta_{t}^{(n)}$ ($|a|\varphi(|a|)\leq\varphi(1)\leq 1$)
and $|\theta|\varphi(|x_{2}(\theta)|)\leq|\mu_{t-1/n}^{(n)*}|+\eta_{t-1/n}^{(n)}$ by analogous arguments. Finally, note that for any fixed $t\geq 1$,
$\eta_{t-1/n}^{(n)}\to\eta_{t}$
since
$\Big{(}\eta_{t-1/n}^{(n)}\Big{)}^{-2}=\Big{(}\eta_{t}^{(n)}\Big{)}^{-2}-n^{-1}\sigma_{t}^{-2}=\eta_{t}^{-2}-n^{-1}\sigma_{t}^{-2}$
and
$\frac{1}{\sqrt{2\pi}\eta_{t}(nt-1)}\to 0$
as $n\to\infty$.
@@
Hence, the time-$t$ strategy $a_{it}=1$ iff $\mu_{it}\geq\mu_{t}^{(n)*}$” is an $\epsilon(n,t)$-equilibrium at time $t$, since we may bound the time-$t$ expected payoff of using this strategy in the contemporaneous game uniformly across players’ time-$t$ information sets $\mu_{it}$: if $\mu_{it}\geq\mu_{t}^{(n)*}$,
@@
Appendix III Signals about past play
@@
Now suppose that the statement holds for t1t-1. By assumption, we have $\Phi^{-1}(\lambda_{t-1})=(\theta-\mu_{t-1}^{*})/\eta_{t-1}$, which implies